OpenStudy (anonymous):

The diagonal of a square has a length of 10 cm. Find the length of a side of the square.

OpenStudy (anonymous):

@manguitar diagonal of square = sqrt(2) * Side of square plug in the value of diagonal and then u will get your side

OpenStudy (anonymous):

could you explain a litttle bit more?

OpenStudy (anonymous):

 Let us assume side of square be 'a' in order to determine d (diagonal) , we used pythagoras theorem \[d^2 = a^2+a^2\] \[d^2 = 2a2\] \[d = \sqrt(2) * a\] where a = side of square and d = diagonal of square